# Characteristic Root Analysis and Regularization for Linear Time Series Forecasting

### v2

@@ Abstract
Time series forecasting remains a critical challenge across numerous domains, yet the effectiveness of complex models often varies unpredictably across datasets. Recent studies highlight the surprising competitiveness of simple linear models, suggesting that their robustness and interpretability warrant deeper theoretical investigation. This paper presents a systematic study of linear models for time series forecasting, with a focus on the role of characteristic roots in temporal dynamics. We begin by analyzing the noise-free setting, where we show that characteristic roots govern long-term behavior and explain how design choices such as instance normalization and channel independence affect model capabilities. We then extend our analysis to the noisy regime, revealing that models tend to produce spurious roots. This leads to the identification of a key data-scaling property: mitigating the influence of noise requires disproportionately large training data, highlighting the need for structural regularization. To address these challenges, we propose two complementary strategies for robust root restructuring. The first uses rank reduction techniques, including Reduced-Rank Regression (RRR) and Direct Weight Rank Reduction (DWRR), to recover the low-dimensional latent dynamics. The second, a novel adaptive method called Root Purge, encourages the model to learn a noise-suppressing null space during training. Extensive experiments on standard benchmarks demonstrate the effectiveness of both approaches, validating our theoretical insights and achieving state-of-the-art results in several settings. Our findings underscore the potential of integrating classical theories for linear systems with modern learning techniques to build robust, interpretable, and data-efficient forecasting models.
@@ 4.2 Root Purge for Dynamic Rank Adjustment
Here, $\mathcal{G}_{\mathbf{W}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{N\times H}$ denotes a linear transformation governed by the learned weight matrix $\mathbf{W}$. The operator $\mathcal{P}:\mathbb{R}^{N\times H}\to\mathbb{R}^{N\times L}$ ensures dimensional consistency by either cropping or zero-padding the output [^3] In practice: if $H<L$, we zero-pad the output to length $L$ and scale $\lambda$ by $L/H$; if $H\geq L$, we crop to the first $L$ columns and keep $\lambda$ unchanged.. The loss consists of two complementary components: (1) The root-seeking term reflects the standard prediction loss commonly used in time series forecasting; (2) The root-purging term serves as a regularizer, encouraging the model to learn a null space that suppresses noise.
@@ 5.2 Main Results
We evaluate our structural regularization approaches for linear models. Table 1 shows the results with rank reduction using RRR and Root Purge in the frequency domain, and full results are given in Appendix E. Experiments are repeated five times, and we report the average performance. The confidence interval for Root Purge is omitted, as run-to-run differences are negligible except for the Exchange dataset (within $\pm 0.002$). Both RRR and Root Purge consistently outperform other baselines. RRR surpasses methods that require extensive hyperparameter tuning, providing simpler rank adjustment without retraining. Meanwhile, Root Purge pushes the performance limits of linear time series forecasting across multiple benchmarks while maintaining simplicity and computational efficiency. These methods are especially effective on smaller datasets, where models relying solely on data scaling tend to underperform. Since we cannot get ground-truth roots for real datasets, We further validate that our methods bring better characteristic roots on synthetic data in Appendix E.13.
5.3 Model Analysis
@@
Effect on Singular Value Spectrum.
Data Scaling & Noise Robustness Property.
Here we use a toy example to illustrate the scaling behavior of different methods under noisy conditions. Specifically, we examine the effects of two key factors: training data volume and noise magnitude. We construct a synthetic signal $\{y_{t}\}_{t\in[0,T]}$ with both trend and periodic components: $y_{t}=\sin(2t)+\cos(5t)+0.5t+\sigma\cdot\varepsilon_{t}$, where $\varepsilon_{t}\sim\mathcal{N}(0,1)$ and observations are taken at increments of 0.01 in $t$. To examine data scaling behavior, we fix the noise level at $\sigma=0.5$ and vary the size of the training dataset. As shown in Figure 4 (left), baseline models struggle under limited data and require substantially more samples to mitigate the impact of noise, revealing a clear data scaling inefficiency. In contrast, our proposed methods, RRR and Root Purge, maintain consistently strong performance across all dataset sizes, with minimal performance drop as data size decreases. To assess noise robustness, we fix the training interval to $t\in[0,200]$ and vary the noise level $\sigma$. Results in Figure 4 (right) show that while baseline models degrade more significantly with increasing noise, RRR and Root Purge remain stable and robust throughout.
@@
The results presented so far are mainly based on the channel independent (CI) strategy, which models all channels using a shared weight matrix. A key advantage of CI is its ability to aggregate noise from all channels. This serves as implicit data augmentation and, by Proposition 1, enhances model robustness. However, the expressivity of CI is inherently limited by its shared parameterization, creating a tradeoff between robustness and modeling capacity. Here, we consider an alternative design: individual channel (INC) modeling, where each channel is assigned its own linear model (see Table 5 for more details of CI and INC). This increases model capacity and allows the system to better adapt to channel-specific dynamics. However, without the benefit of cross-channel noise sharing, INC models are more sensitive to noise and often underperform in raw form, as shown in Table 2. Notably, when combined with our proposed Root Purge strategy, the INC formulation becomes significantly more robust. Root Purge enables each individual model to effectively suppress channel-specific noise and isolate meaningful dynamics. In several cases, this leads to substantial performance gains over CI models. For example, on the ETTm1 dataset, the INC+Root Purge model even outperforms the previous non-linear state-of-the-art PatchTST (Nie et al., 2022) on the 96- and 192-step forecasting tasks.
@@ 6 Conclusions
Detailed descriptions of the metric and experimental setup, dataset we used, and baseline information are provided in Section 5, Appendix E.1, E.2, and E.3. Detailed descriptions, assumptions, and proofs of theories (Facts, Propositions, Remarks, Claims, and Corollaries) can be found in Appendix D. Source code is available in the supplementary material as a zip file.
@@ B.1 Related Works
Low-rank structure is a pervasive modeling assumption for high-dimensional dynamical systems. By positing an effective low-rank interaction matrix, one can capture large-scale behavior in a tractable form. Recent theoretical and empirical studies [Thibeault et al., 2024] probe this hypothesis by analyzing singular value decay in random graphs and real networks, quantifying effective ranks, and linking rapid singular value decay to accurate dimension reductions of nonlinear dynamics. These findings show that low-rank representations enable faithful reduced-order models of recurrent neural networks and other complex systems. Furthermore, rank reduction and low-rank recovery play a central role across applications [Kishore Kumar and Schneider, 2017, Hu et al., 2021]. Beyond classical truncated SVD (hard thresholding), a rich set of algorithms—convex relaxations with the nuclear norm, nonconvex optimization, iteratively reweighted schemes, and modern SVD-free iterative methods—has emerged [Chen et al., 2021, Radhakrishnan et al., 2025]. Recent advances [Krämer, 2025] revisit affine rank minimization and show that asymptotic log-determinant reweighted least squares can recover low-rank solutions under broad conditions, clarifying both computational and statistical trade-offs of low-rank modeling at scale. In the time-series domain, low-rank structure also provides a powerful inductive bias. For example, Agarwal et al. [2018] reformulates univariate time series into structured low-rank Page matrices so that denoising, imputation, and forecasting reduce to classical matrix estimation problems. More recently, Huang et al. [2025] shows that even ultra-lightweight networks with fewer than 0.4k parameters can achieve state-of-the-art long-term forecasting performance by exploiting low-rank periodic structures and orthogonality-constrained representations, drastically reducing model size and computation while maintaining accuracy.
@@ D.6 Rank reduction and noise robustness
  • Result II (Linear Factor Model). When each observation arises from a linear low-rank model $\mathbf{y_{i}}=\mathbf{L}^{\star}\mathbf{x_{i}}+\mathbf{\epsilon_{i}}$ (latent factors $\mathbf{x_{i}}$ and idiosyncratic noise $\mathbf{\epsilon_{i}}$ [^6] This classical linear factor model should not be confused with our linear forecast model; it aims at recovering a low-rank factor structure from independent samples, whereas ours focuses on forecasting future values from past observations., the sample covariance’s top-$\rho$ eigenspace estimates the population principal subspace with an error that decreases with sample size $n$; projecting observations onto this estimated low-rank subspace reduces variance of the fitted signals and yields improved reconstruction.
@@
$\|\mathbf{\widehat{M}}-\mathbf{M}^{\star}\|_{2}\leq 10\,\sigma\sqrt{n},\quad\|\mathbf{\widehat{M}}-\mathbf{M}^{\star}\|_{F}\leq 10\,\sigma\sqrt{2rn}.$
@@
Step 3: Controlling the subspace estimation error. Since the target matrix $\mathbf{M}^{\star}$ has rank $\rho$ with eigenvalue gap $|\lambda_{\rho}^{\star}|$, the Davis–Kahan theorem [Davis and Kahan, 1970] gives
@@
Under the signal-to-noise condition $\|\mathbf{E}\|_{2}\leq(1-1/\sqrt{2})|\lambda_{\rho}^{\star}|$, this becomes
$\mathrm{dist}(\mathbf{U},\mathbf{U}^{\star})\;\lesssim\;\frac{\sigma\sqrt{n}}{|\lambda_{\rho}^{\star}|}.$
Step 4: Bounds on eigenvalues and low-rank reconstruction. Weyl’s inequality [Weyl, 1912] implies that all spurious eigenvalues beyond rank $r$ are bounded by $\|\mathbf{E}\|_{2}$. For the low-rank reconstruction $\widehat{\mathbf{M}}:=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, the triangle inequality yields
@@
Since the difference has rank at most $2r$, we further obtain
@@ D.7 Analysis of Root Purge
The simultaneous diagonalizability of $\mathbf{X}^{\top}\mathbf{X}$ and $\mathbf{R}^{\top}\mathbf{R}$ should be interpreted as an analytical choice of coordinate system rather than a restrictive modeling assumption. Indeed, the result characterizes shrinkage behavior in the common eigenbasis of the data covariance and the residual covariance. In particular, directions associated with large residual energy experience stronger suppression, while directions well explained by the data are preserved. As the regularization parameter $\lambda$ increases, this selective attenuation progressively eliminates low-energy singular modes, effectively reducing the rank of the solution. This theoretical behavior is clearly reflected in our empirical results, as shown in Figure  3 and Figure 14, where increasing $\lambda$ progressively attenuates smaller singular values. This supports the interpretation that Root Purge acts similarly to rank reduction techniques by selectively suppressing spurious or noisy components (roots) in the solution. Hence, Root Purge can be seen as a principled mechanism for regularizing overparameterized models by promoting low-rank structure.
@@
As a consequence of this proposition, we see that gradient descent reaches an $\epsilon$-stationary point ($\|\nabla J(\mathbf{W})\|_{F}^{2}\leq\epsilon$) in $\mathcal{O}(1/\epsilon)$ iterations. While the IRLS algorithm provides faster convergence through its second-order approximation, it requires solving a linear system with an additional computational cost of $\mathcal{O}(L^{3})$ per iteration due to the matrix inversion of $\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{R}_{k}^{\top}\mathbf{R}_{k}$ (computational costs in matrix multiplications are not counted as it is also required for other algorithms). In contrast, gradient descent avoids this cubic scaling, making it preferable for extended lookback window problems where $L\gg 1$. Note that this convergence result applies to any objective function satisfying the Lipschitz gradient condition. In particular, it applies to our specific optimization problem:

### v3

@@ Abstract
Time series forecasting remains a critical challenge across numerous domains, yet the effectiveness of complex models often varies unpredictably across datasets. Recent studies highlight the surprising competitiveness of simple linear models, suggesting that their robustness and interpretability warrant deeper theoretical investigation. This paper presents a systematic study of linear models for time series forecasting, with a focus on the role of characteristic roots in temporal dynamics. We begin by analyzing the noise-free setting, where we show that characteristic roots govern long-term behavior and explain how design choices such as instance normalization and channel independence affect model capabilities. We then extend our analysis to the noisy regime, revealing that models tend to produce spurious roots. This leads to the identification of a key data-scaling property: mitigating the influence of noise requires disproportionately large training data, highlighting the need for structural regularization. To address these challenges, we propose two complementary strategies for robust root restructuring. The first uses rank reduction techniques, including Reduced-Rank Regression (RRR) and Direct Weight Rank Reduction (DWRR), to recover the low-dimensional latent dynamics. The second, a novel adaptive method called Root Purge, encourages the model to learn a noise-suppressing null space during training. Extensive experiments on standard benchmarks demonstrate the effectiveness of both approaches, validating our theoretical insights and achieving state-of-the-art results in several settings. Our findings underscore the potential of integrating classical theories for linear systems with modern learning techniques to build robust, interpretable, and data-efficient forecasting models. The code is publicly available at: https://github.com/Wangzzzzzzzz/Root-Purge-for-Time-Series-Forecasting.
@@ 4.1 Rank Reduction Methods
See Appendix D.5 for additional theoretical analysis of Rank-Reduction methods.
@@ 4.2 Root Purge for Dynamic Rank Adjustment
Here, $\mathcal{G}_{\mathbf{W}}:\mathbb{R}^{N\times L}\to\mathbb{R}^{N\times H}$ denotes a linear transformation governed by the learned weight matrix $\mathbf{W}$. The operator $\mathcal{P}:\mathbb{R}^{N\times H}\to\mathbb{R}^{N\times L}$ ensures dimensional consistency by either cropping or zero-padding the output [^3] In practice: if $H<L$, we zero-pad the output to length $L$ and scale $\lambda$ by $L/H$; if $H\geq L$, we crop to the first $L$ columns and keep $\lambda$ unchanged. [^4] For more memory-efficient training, we apply a stop-gradient operation to $\mathcal{P}(\cdot)$. Empirically, this has minimal impact on the performance of Root Purge. See Appendix E.14 for further details.. The loss consists of two complementary components: (1) The root-seeking term reflects the standard prediction loss commonly used in time series forecasting; (2) The root-purging term serves as a regularizer, encouraging the model to learn a null space that suppresses noise.
@@
To see additional theoretical analysis of Root Purge, please refer to Appendix D.7.
@@ 5.2 Main Results
We evaluate our structural regularization approaches for linear models. Table 1 shows the results with rank reduction using RRR and Root Purge in the frequency domain, and results on time domain models are given in Appendix E.7 and E.8. Experiments are repeated five times, and we report the average performance. The confidence interval for Root Purge is omitted, as run-to-run differences are negligible except for the Exchange dataset (within $\pm 0.002$). Both RRR and Root Purge consistently outperform other baselines. RRR surpasses methods that require extensive hyperparameter tuning, providing simpler rank adjustment without retraining. Meanwhile, Root Purge pushes the performance limits of linear time series forecasting across multiple benchmarks while maintaining simplicity and computational efficiency. These methods are especially effective on smaller datasets, where models relying solely on data scaling tend to underperform, in line with our theoretical analysis.
5.3 Theoretical Properties Verification
In this section, we validate the theoretical properties of our proposed methods in three stages: from singular value shrinkage (the mathematical mechanism) to improved data scaling (the key practical challenge) and finally back to characteristic root restructuring (the foundational motivation). We then further assess their sensitivity to hyperparameters and the possibility of alternative architectural choices in the next section.
Effect on Singular Value Spectrum.
Data Scaling & Noise Robustness Property.
We further use a toy example to illustrate the scaling behavior of different methods under noisy conditions. Specifically, we examine the effects of two key factors: training data volume and noise magnitude. We construct a synthetic signal $\{y_{t}\}_{t\in[0,T]}$ with both trend and periodic components: $y_{t}=\sin(2t)+\cos(5t)+0.5t+\sigma\cdot\varepsilon_{t}$, where $\varepsilon_{t}\sim\mathcal{N}(0,1)$ and observations are taken at increments of 0.01 in $t$. To examine data scaling behavior, we fix the noise level at $\sigma=0.5$ and vary the size of the training dataset. As shown in Figure 3 (left), baseline models struggle under limited data and require substantially more samples to mitigate the impact of noise, revealing a clear data scaling inefficiency. In contrast, our proposed methods, RRR and Root Purge, maintain consistently strong performance across all dataset sizes, with minimal performance drop as data size decreases. To assess noise robustness, we fix the training interval to $t\in[0,200]$ and vary the noise level $\sigma$. Results in Figure 3 (right) show that while baseline models degrade more significantly with increasing noise, RRR and Root Purge remain stable and robust throughout.
Table 2: Comparison of Root Distance to Ground Truth across models (root pairings determined by Hungarian algorithm (Kuhn, 1955)). The results demonstrate the effectiveness of Rank Reduction and Root Purge in recovering roots closer to the true values compared to standard linear models.
Model Root Distance to Ground Truth (mean $\pm$ std)
RRR $\mathbf{0.036\pm 0.014}$
Root Purge $\mathbf{0.045\pm 0.009}$
Standard Linear Model (OLS) $0.064\pm 0.025$
Analysis of Characteristic Roots.
Finally, to close the loop of on our framework, we hereby directly evaluate how Rank Reduction Regression (RRR) and Root Purge affect the learned root distribution in a controlled synthetic setting. Such a direct comparison is usually infeasible on real-world data because the underlying noise-free dynamics and their true characteristic roots are unknown. Therefore, we conduct experiments using the aforementioned synthetic dynamics: $y_{t}=\sin(2t)+\cos(5t)+0.5t+\sigma\cdot\varepsilon_{t}$. For each method, we compute the mean and standard deviation of the distances between all learned roots and the ground truth roots, with root pairs matched using the Hungarian algorithm (Kuhn, 1955). As shown in Table 2, the roots estimated by OLS tend to drift farther from the ground-truth characteristic roots, whereas those estimated by Root Purge and RRR remain closer. This observation confirms that our proposed techniques indeed help prevent the learning of spurious roots, thereby leading to the performance improvements demonstrated earlier. Additional experimental details and qualitative evaluations are also provided in Appendix E.13.
5.4 Practical Robustness Analysis
@@
The results presented so far are mainly based on the channel independent (CI) strategy, which models all channels using a shared weight matrix. A key advantage of CI is its ability to aggregate noise from all channels. This serves as implicit data augmentation and, by Proposition 1, enhances model robustness. However, the expressivity of CI is inherently limited by its shared parameterization, creating a tradeoff between robustness and modeling capacity. Here, we consider an alternative design: individual channel (INC) modeling, where each channel is assigned its own linear model (see Table 6 for more details of CI and INC). This increases model capacity and allows the system to better adapt to channel-specific dynamics. However, without the benefit of cross-channel noise sharing, INC models are more sensitive to noise and often underperform in raw form, as shown in Table 3. Notably, when combined with our proposed Root Purge strategy, the INC formulation becomes significantly more robust. Root Purge enables each individual model to effectively suppress channel-specific noise and isolate meaningful dynamics. In several cases, this leads to substantial performance gains over CI models. For example, on the ETTm1 dataset, the INC+Root Purge model even outperforms the previous non-linear state-of-the-art PatchTST (Nie et al., 2022) on the 96- and 192-step forecasting tasks. These results highlight how our method unlocks alternative, more expressive architectures that would otherwise be infeasible.
@@ 6 Conclusions
Detailed descriptions of the metric and experimental setup, dataset we used, and baseline information are provided in Section 5, Appendix E.1, E.2, and E.3. Detailed descriptions, assumptions, and proofs of theories (Facts, Propositions, Remarks, Claims, and Corollaries) can be found in Appendix D. Source code is available in the supplementary material as a zip file. You can also find our source code from https://github.com/Wangzzzzzzzz/Root-Purge-for-Time-Series-Forecasting.
@@ B.1 Related Works
Low-rank structure is a pervasive modeling assumption for high-dimensional dynamical systems Agarwal et al. [2018], Liu [2022]. By positing an effective low-rank interaction matrix, one can capture large-scale behavior in a tractable form. Recent theoretical and empirical studies [Thibeault et al., 2024] probe this hypothesis by analyzing singular value decay in random graphs and real networks, quantifying effective ranks, and linking rapid singular value decay to accurate dimension reductions of nonlinear dynamics. These findings show that low-rank representations enable faithful reduced-order models of recurrent neural networks and other complex systems. Furthermore, rank reduction and low-rank recovery play a central role across applications [Kishore Kumar and Schneider, 2017, Hu et al., 2021]. Beyond classical truncated SVD (hard thresholding), a rich set of algorithms—convex relaxations with the nuclear norm, nonconvex optimization, iteratively reweighted schemes, and modern SVD-free iterative methods—has emerged [Chen et al., 2021, Radhakrishnan et al., 2025]. Recent advances [Krämer, 2025] revisit affine rank minimization and show that asymptotic log-determinant reweighted least squares can recover low-rank solutions under broad conditions, clarifying both computational and statistical trade-offs of low-rank modeling at scale. In the time-series domain, low-rank structure also provides a powerful inductive bias. For example, Agarwal et al. [2018] reformulates univariate time series into structured low-rank Page matrices so that denoising, imputation, and forecasting reduce to classical matrix estimation problems. More recently, Huang et al. [2025] shows that even ultra-lightweight networks with fewer than 0.4k parameters can achieve state-of-the-art long-term forecasting performance by exploiting low-rank periodic structures and orthogonality-constrained representations, drastically reducing model size and computation while maintaining accuracy.
@@ D.6 Rank reduction and noise robustness
  • Result II (Linear Factor Model). When each observation arises from a linear low-rank model $\mathbf{y_{i}}=\mathbf{L}^{\star}\mathbf{x_{i}}+\mathbf{\epsilon_{i}}$ (latent factors $\mathbf{x_{i}}$ and idiosyncratic noise $\mathbf{\epsilon_{i}}$) [^7] This classical linear factor model should not be confused with our linear forecast model; it aims at recovering a low-rank factor structure from independent samples, whereas ours focuses on forecasting future values from past observations., the sample covariance’s top-$\rho$ eigenspace estimates the population principal subspace with an error that decreases with sample size $n$; projecting observations onto this estimated low-rank subspace reduces variance of the fitted signals and yields improved reconstruction.
@@
$\|\mathbf{\widehat{M}}-\mathbf{M}^{\star}\|_{2}\leq 10\,\sigma\sqrt{n},\quad\|\mathbf{\widehat{M}}-\mathbf{M}^{\star}\|_{F}\leq 10\,\sigma\sqrt{2\rho n}.$
@@
Step 3: Controlling the subspace estimation error. Since the target matrix $\mathbf{M}^{\star}$ has rank $\rho$ with eigenvalue gap $|\lambda_{\rho}^{\star}|$, then under the signal-to-noise condition $\|\mathbf{E}\|_{2}\leq(1-1/\sqrt{2})|\lambda_{\rho}^{\star}|$, the Davis–Kahan theorem [Davis and Kahan, 1970] gives
@@
Step 4: Bounds on eigenvalues and low-rank reconstruction. Weyl’s inequality [Weyl, 1912] implies that all spurious eigenvalues beyond rank $\rho$ are bounded by $\|\mathbf{E}\|_{2}$. For the low-rank reconstruction $\widehat{\mathbf{M}}:=\mathbf{U}\mathbf{\Lambda}\mathbf{U}^{\top}$, the triangle inequality yields
@@
Since the difference has rank at most $2\rho$, we further obtain
@@ D.7 Analysis of Root Purge
The simultaneous diagonalizability of $\mathbf{X}^{\top}\mathbf{X}$ and $\mathbf{R}^{\top}\mathbf{R}$ should be interpreted as an analytical choice of coordinate system rather than a restrictive modeling assumption. Indeed, the result characterizes shrinkage behavior in the common eigenbasis of the data covariance and the residual covariance. In particular, directions associated with large residual energy experience stronger suppression, while directions well explained by the data are preserved. As the regularization parameter $\lambda$ increases, this selective attenuation progressively eliminates low-energy singular modes, effectively reducing the rank of the solution. This theoretical behavior is clearly reflected in our empirical results, as shown in Figure  2 and Figure 14, where increasing $\lambda$ progressively attenuates smaller singular values. This supports the interpretation that Root Purge acts similarly to rank reduction techniques by selectively suppressing spurious or noisy components (roots) in the solution. Hence, Root Purge can be seen as a principled mechanism for regularizing overparameterized models by promoting low-rank structure.
@@
As a consequence of this proposition, we see that gradient descent reaches an $\epsilon$-stationary point ($\|\nabla J(\mathbf{W})\|_{F}^{2}\leq\epsilon$) in $\mathcal{O}(1/\epsilon)$ iterations. While recursively solving least square may provides faster convergence through its second-order approximation, it requires solving a linear system with an additional computational cost of $\mathcal{O}(L^{3})$ per iteration due to the matrix inversion of $\mathbf{X}^{\top}\mathbf{X}+\lambda\mathbf{R}_{k}^{\top}\mathbf{R}_{k}$ (computational costs in matrix multiplications are not counted as it is also required for other algorithms). In contrast, gradient descent avoids this cubic scaling, making it preferable for extended lookback window problems where $L\gg 1$. Note that this convergence result applies to any objective function satisfying the Lipschitz gradient condition. In particular, it applies to our specific optimization problem: